# Fundamental Oscillations of Massive Boson Stars - II

### v1

@@ Abstract
Boson Stars (BSs) are macroscopic self-gravitating configurations made of complex scalar fields. These exotic compact objects (ECO) would manifest as dark Boson stars and can contribute to a certain fraction of the dark matter (DM) in the universe. In this work, we study the fundamental non-radial oscillations ($f$-modes) of massive BSs and the associated gravitational wave (GW) emission. We consider massive scalar BSs having the potential of the form $V(\phi)=\frac{1}{2}m^{2}|\phi|^{2}+\frac{1}{4}\lambda|\phi|^{4}$, restricting to the strong-coupling regime ($\lambda\gg m^{2}/M_{Pl}^{2}$) where solutions resembling fermionic stars are known to exist. We first review the available parameter space for scalar DM that can form massive BSs and enlist various constraints. We fit and provide simple analytical relations connecting various macroscopic observables for BSs, which can be directly incorporated into future studies of massive BSs throughout the strong coupling regime without requiring any numerical computation. We then solve for the non-radial $l=2$ fundamental quasinormal modes ($f$-modes) for massive BSs. Scaling relations for $f$-mode equations have been reported in another work. Using these, we perform a complete study of these oscillations spanning the entire available parameter space of massive BSs and provide analytical fits for the $f$-mode characteristics. We further study the universal relations for BSs and reveal the parameter space that is sensitive to the current and future planned GW detectors. We find that different parts of the parameters space can, in principle, be probed by the LISA, LIGO, and NEMO detectors. Finally, we briefly discuss the detectability of $f$-modes from BSs that have not been explored before.
@@ 1 Introduction
Boson stars (BSs) are stellar configurations of bosonic particles (see [1, 2] for updated review on BSs). The formation channels for BSs include primordial density fluctuations in the scalar field, gravothermal collapse in an early matter-dominated era of the universe [3], or a process known as gravitational cooling [4]. These are some of the possible exotic compact objects (ECOs) that can mimic other compact objects like black holes (BHs) acting as BH mimickers, as well as contribute to the total DM budget (see [5] for a review on ECOs). When a real scalar field describes these stars, they form a Soliton star, also called Oscillaton [6]. In these stars, the boson number is not conserved. When the field is complex, the stellar configurations formed thereby are called Boson Stars [6], the boson number of which is conserved. When a particle is modelled using vector boson (spin-1), the stars obtained are called Proca Stars.
The study of such stars started with the self-gravitating configurations of the electromagnetic field [7]. This was then applied to massive-scale fields by solving the Einstein-Klein-Gordon equations [8, 9, 10] and extended later to include self-interactions [11, 12]. Various types of BSs have been studied in the literature. Based on the interaction potential, they are classified into different types of BSs like mini-BS (no interaction), massive BS, axion star, solitonic star, etc. In this work, we will focus on the second type, the massive BSs, which is the simplest form of potential consisting of quartic self-interaction capable of forming massive compact objects as shown in [12].
@@
Being a candidate for dark matter, these particles are assumed to couple sufficiently weakly to electromagnetic (EM) photons and other SM particles. This means that non-accreting BSs are not observable via conventional EM telescopes. This situation is similar to that of black holes (BHs), making them a good candidate for BH mimickers. In the case of accreting systems, the BSs can be distinguished using the EM observations based on the analysis of shadow region [13] and line broadening [14]. These methods are inaccessible in the case of non-accreting BSs. However, the successful detection of gravitational waves [15] has opened up a new window to probe the dark sectors of the universe. There have been many works on this front in the last few decades. These, along with search results so far, have been briefly summarized below.
Head-on collisions of BSs and the associated GW emission were studied in [16, 17, 18, 19]. Gravitational echoes from ECOs were explored in [20]. Recently, [21] performed the first simulation of a binary merger of massive BSs and [22] for solitonic stars.
GW190521 was a BBH merger event with at least one of the components having mass lying in the black hole mass gap [23]. There were several beyond standard model (BSM) explanations for explaining this merger [24] along with the possibility of BS mergers [25], in particular, head-on collision of high-spin Proca stars. This was the first instance of the possibility of having detected BS as a BH mimicker through a real GW event, however, later analysis of higher order modes [26] of the ringdown phase ruled out this possibility in favour of BH.
Quasi-normal oscillation modes of stars are important in the context of binary pre- and post-merger as well as isolated system excitations. [27] was the first work to explore the quasinormal modes of BSs. There have been a few studies on BS QNMs following that [28, 29, 30, 31, 32, 33].  [31] is the first work to study the $f$-modes for massive BSs in the strong-interaction limit. However, the work was restricted to stellar mass BSs and select model parameters, as will be discussed later. Another work [33] appeared during the completion of this work that explored BS $f$-modes for select parameter sets. We report scaling relations in a separate publication [34] and therein compare the BS QNMs with those of BHs and neutron stars (NSs). Spinning BSs are known to form toroidal configurations instead of spherical [35]. We restrict to the non-spinning case in this work.
We extend the work of [31, 33] to make a more general case. As will be discussed, a large parameter space in terms of DM mass and its self-interaction strength $\lambda$ is available for BSs. In [34], we show for the first time that scaling relations are followed for the $f$-mode frequency and damping time in the strong-interaction limit. This scaling result is significant, as it allows us to comment on $f$-mode characteristics of configurations of arbitrary parameters within the strong-interaction limit ($\Lambda_{int}\gg 1$). We provide analytical fits to the static as well as $f$-mode solutions for massive BSs and also study the universal relations. We estimate the parameter space accessible for existing and future GW detectors and comment on the detectability of BSs using $f$-modes.
@@
The article is organized as follows: In Sec. 2 we provide the details of the BS model, describing the scaling property of massive BS. We discuss the available scalar DM parameter space in Sec. 3, reviewing all the available constraints. We present all the results in Sec 4, discussing the static observables, $f$-modes, and the corresponding universal relations. We also comment on the parameter space observable using GW from $f$-modes from massive BSs and discuss the detectability by various GW detectors. We summarize our results in Sec. 5, providing a comparison with earlier works and outlining the future scope.
We use the convention $\hbar=c=1$ and hence $G=1/M_{Pl}^{2}$ throughout this work.
@@ 2 Model
We describe DM by scalar field self-interacting via quartic self-interactions as elaborated in [12, 31]. The interaction strength is given by $\lambda$.
@@
Here $\lambda$ is the self-interaction coupling, and $m$ is the mass of the DM particle. A dimensionless parameter is defined as
$\Lambda=\frac{\lambda M_{Pl}^{2}}{4\pi m^{2}}~{}.$ (2.3)
For $\Lambda\gg 1$[12] showed that we could form massive stars whose mass resembles that of a degenerate fermionic star. In this case, $M_{max}=0.22\sqrt{\Lambda}M_{Pl}^{2}/m=0.06\sqrt{\lambda}M_{pl}^{3}/m^{2}$ [12]. The macroscopic properties are dictated by the effective factor defined as [21]
@@
Using this, we define the dimensionless mass ($M^{\prime}$) and dimensionless radius ($R^{\prime}$) as $M^{\prime}=M/(xM_{Pl}^{3})~{},R^{\prime}=R/(xM_{Pl})$. The dimensionless energy density and pressure are scaled as $\rho^{\prime}=\rho x^{2}~{},p^{\prime}=px^{2}$.
We consider here the case of self-interacting massive BSs, for which $\Lambda\gg 1$ and the scalings hold. We restrict our analysis to $\Lambda>1000$, referred to as the strong-interaction limit [36]. In this limit, the Einstein-Klein-Gordon system of equations for the scalar field resembles that of a perfect fluid star with the effective equation of state given by [12, 37, 21],
@@
For $\rho^{\prime}\ll 1$, $p^{\prime}=\rho^{\prime 2}/4$, i.e. we get a polytrope EoS with polytropic index $n=1$. For $\rho^{\prime}\gg 1$, we get an ultra-relativistic EoS $p^{\prime}=\rho^{\prime}/3$. We will show later that this condition is not reached inside BSs (Appendix. B).
For the $f$-mode frequencies and damping times, we solve the $f$-mode eigenfrequencies using full-GR formalism and use the following scaling relations as reported in [34]
$f$ $=f^{\prime}/(xM_{Pl})$ (2.6)
$\tau$ $=\tau^{\prime}(xM_{Pl})~{}.$ (2.7)
The formalism and derivation of scaling have been provided in detail in the Appendix of [34].
@@ 3 Parameter Space
Fig. 1 shows the available parameter space in the $\lambda-m$ space for scalar DM. For mass, we show the range $10^{-25}\text{eV}<m<100\text{GeV}$ covering the mass of fuzzy DM or ultralight ALPs to roughly the mass of WIMPs. We consider $\lambda$ upto $\lambda=15$. This encompasses the upper limit of $\sim 4\pi$ as imposed by [38]. All the constraints are discussed in detail in the next section. The values $\lambda\lesssim 10^{-100}$ are irrelevant as lower values consistent with observations do not fall in the strong-interaction limit considered here. The yellow lines denote contours of fixed $\Lambda$, and this analysis is only applicable to $\Lambda>1000$, i.e., the strong-interaction regime. The translucent region is the Kaup limit where $\Lambda\ll 1$ and interactions can be ignored. The gray lines show contours where $\sigma/m$ is constant, and the region between them is the region permitted by other astrophysical observations. We note that this is a very narrow region in the parameter space; however, the lower bound (dashed gray line) is not strict. The red line marks the constraint from CMB and LSS data [39]. The region above this line is excluded. The blue region is excluded by the recently set lower limit in scalar DM mass [40] using dwarf galaxies.
@@
The maximum mass of the Bosonic dark star is shown in colour. This can be calculated for a given set of parameters $(\lambda,m)$ is given by [12]: $M_{max}=0.22\sqrt{\Lambda}\frac{M_{Pl}^{2}}{m}=0.06\frac{\sqrt{\lambda}}{m^{2}}M_{Pl}^{3}=0.06xM_{Pl}^{3}$. The colour scale saturates into yellow at $M_{max}=20^{22}M_{\odot}$, which is the rough estimate for the universe’s mass. We plot the contours of some fixed mass values. We find that the $M_{max}$ ranges a larger region in mass spanning mountain mass to galactic mass (see. caption of Fig. 1). The arrows indicate the regions allowed by various constraints. The only region bounded by the gray ($\sigma/m$), red (CMB, LSS), and blue (dwarf galaxies) lines is the allowed region for scalar DM and, hence, BSs. Further, this work is only applicable in the region bounded by the yellow (strong-interaction regime) line, i.e., the strong interaction regime.
@@
Flores et al. [31] restricted their study to stellar mass BSs and used selected values of $m\in\{1,1.25,1.5,1.75,2\}m_{n}$ and $a\in\{5,10,15,20\}$ fm to study the $f$-modes. The scattering length parameter $a=\lambda/8\pi m$ was used to fix interaction strength instead of $\lambda$ in [31]. These are marked by a purple patch in Fig. 1. Maselli et al. [41] studied the $I$-Love-$Q$ relations using a select set of model values. The parameters chosen we $m\in\{300,400\}$ MeV and $\lambda\in\{0,0.5,1.5\}\pi$. Celato et al. [33] explored the $f$-mode universal relations for the same parameters.  These are marked in orange in Fig. 1. It is clear that both these patches lie near the $M_{max}\sim 1M_{\odot}$ contour and focus on the configurations in the NS mass range.
@@
Figure 1: Figure shows the available parameter space in the $\lambda-m$ space. The color indicates the logarithm of the maximum mass of a stable BS for a given pair of ($\lambda$,m). We show select contour lines (black) to get an idea of masses: $\log_{10}M=$18 (mountain mass; this is also the lower bound on primordial BH set by Hawking radiation as lighter PBHs would evaporate within Hubble time); -9 (asteroid belt mass); -6 (Earth-like planet mass); -3 (giant planet mass); 0 (stellar mass); 2 (giant massive star mass); 6 and 9 (SMBH mass); 12 (galaxy mass). The red line gives an upper limit $\lambda$ and a lower limit on $m$ from CMB and LSS data [39]. The blue line marks the lower limit on scalar DM mass [40] from dwarf galaxies. The yellow contours represent values of $\Lambda$. We restrict to the strong-interaction limit $\Lambda>1000$ in this work. $\Lambda\ll 1$ is the non-interacting/Kaup limit shaded in gray. The bisque band is the from the bounds on $\sigma/m$. The upper limit and lower limit are marked by a solid and dashed line, respectively. The lower bound is not strict. See text for details. The purple patch denotes the parameters used in [31] to study $f$-modes of massive BS. The orange patch indicates the parameters used by [41, 33] to study the $I$-Love-$Q$ and $f$-mode universal relations.

### v3

# Detectability of Massive Boson Stars using Gravitational Waves from Fundamental Oscillations

@@ Abstract
Boson Stars are macroscopic self-gravitating configurations made of complex scalar fields. These exotic compact objects would manifest as dark Boson stars and, in the absence of electromagnetic signatures, could mimic properties of compact stars in the gravitational wave spectrum. In a recent study, using the simplest potential for massive Boson stars, we demonstrated that fundamental non-radial oscillations ($f$-modes) obey scaling relations that allow them to be distinguished from neutron stars and black holes. In this work, we provide analytical fits for these scaling relations, valid for the dark matter parameter space compatible with current astrophysical and cosmological data, that can be directly incorporated into future studies of massive Boson stars in the strong coupling regime, avoiding the need for numerical calculations. We also provide analytical fits for empirical and universal relations for gravitational wave asteroseismology, which can be used to infer microscopic dark matter properties following a successful detection. Further, we investigate the possibility of detection of $f$-modes and the dark matter parameter space that can be probed with current and future gravitational wave detectors across multiple frequency bands. Assuming a burst GW model and demanding a signal-to-noise ratio of 5, we show that the current and future detectors can, in principle, probe Boson star $f$-modes up to cosmological distances: 1 Mpc with aLIGO, 30 Mpc with Cosmic Explorer and Einstein Telescope, and in the best case scenario, about 300 Mpc with LISA.
@@ 1 Introduction
Boson stars (BSs) are stellar configurations of bosonic particles (see [Visinelli2021, LieblingPalenzuela2023] for an updated review on BSs). The formation channels for BSs include primordial density fluctuations in the scalar field [LiddleMadsen1992], gravothermal collapse in an early matter-dominated era of the universe [Ralegankar2024], or a process known as gravitational cooling [Seidel1994]. These are some of the possible exotic compact objects (ECOs) that can mimic other compact objects like neutron stars (NSs) and black holes (BHs), acting as BH mimickers, as well as contribute to the total dark matter (DM) budget (see [CardosoPani2019] for a review on ECOs). Various types of BSs have been studied in the literature [Wheeler1955, Feinblum1968, Kaup1968, Ruffini1969, Mielke1981, Colpi1986]. Based on the interaction potential, they are classified into different types of BSs like mini-BS (no interaction), massive BS, axion star, solitonic star, etc. In this work, we will focus on massive BSs, with the simplest form of potential consisting of quartic self-interaction capable of forming massive compact objects as shown in [Colpi1986].
@@
Being a candidate for dark matter, these particles are assumed to couple sufficiently weakly to electromagnetic (EM) photons and other Standard Model particles. This means that non-accreting BSs are not observable via conventional EM telescopes. Further, massive BSs are compact with compactness $C\lesssim 0.16$ and can have a wide range of mass depending on the model parameters, overlapping the range of BH mass. This makes them a good candidate for BH mimickers. In the case of accreting systems, the BSs can be distinguished using the EM observations based on the analysis of the shadow region [Olivares2020, Rosa2022, Rosa2023] and line broadening [Rosa2024]. These methods are inaccessible in the case of non-accreting BSs. However, the successful detection of gravitational waves (GWs) [GWDetection2016] has opened up a new window to probe the dark sectors of the universe. There have been many works and searches for BSs using GWs in the last few decades [Palenzuela2007, Palenzuela2008, Cardoso2016, Palenzuela2017, Urbano2019, Abbott2020, Sakstein2020, Bustillo2021, Capano2023, Cipriani2024, Evstafyeva2024].
Quasinormal oscillation modes of stars are important in the context of binary pre- and post-merger as well as isolated system excitations. [Kojima1991] was the first work to explore the quasinormal modes (QNMs) of BSs. There have been a few studies on BS QNMs following that [Yoshida1994, Balakrishna1998, Macedo2013a, Flores2019, Kain2021, Celato2025]. $f$-modes, the main motivation of this work, are fundamental (zero nodes) non-radial oscillation modes of a star. Typically, oscillation modes of compact stars like NSs are classified based on the restoring force. In the case of $f$-modes, the restoring force is the fluid pressure. Massive BSs in the strong interaction limit have an equation of state (EoS) equivalent to a polytropic fluid, as will be shown in Sec. 2. This allows us to study the BS $f$-modes, where the pressure is sourced by the repulsive self-interaction.
[Flores2019] is the first work to study the $f$-modes for massive BSs in the strong-interaction limit. Another work [Celato2025] appeared during the completion of this work that explored BS $f$-modes. However, these studies were restricted to select model parameters. In a separate publication [shirke_BS_fmodes_short], we showed for the first time that scaling relations are obeyed by the $f$-mode frequency and damping time in the strong-interaction limit. This scaling result is significant, as it allows us to comment on $f$-mode characteristics of configurations of arbitrary parameters within the strong-interaction limit ($\Lambda\gg 1$; defined in Eq. 2.3) and can be directly used for studies of massive BSs. In this work, we apply these scaling relations independent of microscopic parameters to obtain analytical fits connecting macroscopic structure properties and $f$-mode characteristics of massive BSs, avoiding the need for numerical calculations and reducing computation costs. We also provide universal relations that can be used to infer stellar properties from GW observations and investigate the prospects of detecting GWs from $f$-modes from massive BSs using the current and future detectors.
@@
The article is organized as follows: In Sec. 2 we provide the details of the BS model, describing the scaling property of massive BS. We discuss the available scalar DM parameter space in Sec. 3, reviewing all the available constraints. We present all the results in Sec 4, discussing the static observables, $f$-modes, and the corresponding universal relations. We also comment on the parameter space observable using GWs from $f$-modes from massive BSs and discuss the detectability by various GW detectors. We summarize our results in Sec. 5, providing a comparison with earlier works and outlining the future scope. We use the convention $\hbar=c=1$ and hence $G=1/M_{\rm Pl}^{2}$ throughout this work.
@@ 2 Model
We describe DM by a scalar field self-interacting via quartic self-interactions as elaborated in [Colpi1986, Flores2019]. The interaction strength is given by $\lambda$.
@@
Here $\lambda$ is the self-interaction coupling and $m$ is the mass of the DM particle. A dimensionless parameter is defined as
$\Lambda=\frac{\lambda M_{\rm Pl}^{2}}{4\pi m^{2}}~.$ (2.3)
For $\Lambda\gg 1$[Colpi1986] showed that we could form massive stars whose mass resembles that of a degenerate fermionic star. In this case, $M_{\rm max}=0.22\sqrt{\Lambda}M_{\rm Pl}^{2}/m=0.06\sqrt{\lambda}M_{\rm Pl}^{3}/m^{2}$ [Colpi1986]. The macroscopic properties are dictated by the effective factor defined as [Cipriani2024]
@@
Using this, we define the dimensionless mass ($M^{\prime}$) and dimensionless radius ($R^{\prime}$) as $M^{\prime}=M/(xM_{\rm Pl}^{3})~,R^{\prime}=R/(xM_{\rm Pl})$. The dimensionless energy density and pressure are scaled as $\rho^{\prime}=\rho x^{2}~,p^{\prime}=px^{2}$. The dimensionless compactness $C=GM/R=M/(M_{\rm Pl}^{2}R)=M^{\prime}/R^{\prime}=C^{\prime}$ remains unaltered. Thus, $C=C^{\prime}$ can be used interchangeably.
We consider here this case of self-interacting massive BSs with $\Lambda\gg 1$, for which the scalings hold. We restrict our analysis to $\Lambda>1000$, referred to as the strong-interaction limit [Seoane2010]. All the results in this work are applicable only in this strong-interaction regime. This has been marked by the yellow arrow in Fig. 1. In this limit, the Einstein-Klein-Gordon system of equations for the scalar field resembles that of a perfect fluid star with the effective equation of state given by [Colpi1986, Karkevandi2022, Cipriani2024],
@@
For $\rho^{\prime}\ll 1$, $p^{\prime}=\rho^{\prime 2}/4$, i.e., we get a polytrope EoS with polytropic index $n=1$. For $\rho^{\prime}\gg 1$, we get an ultra-relativistic EoS $p^{\prime}=\rho^{\prime}/3$.
The TOV equations, when written in terms of scaled mass ($M^{\prime}$) and radius ($R^{\prime}$), are known to follow self-similarity, i.e., the equations become independent of model parameter $x$ (see appendix of Maselli et al. [Maselli2017]). This results in a unique mass-radius relation $M^{\prime}-R^{\prime}$, which can be used to obtain the MRM-R curve for any parameters using the scaling relations. The shape of the solution remains the same. Upon solving these scaled TOV equations, we find that the maximum central density reached for a maximum TOV mass configuration is $\rho^{\prime}=1.6$. Thus, $\rho^{\prime}\gg 1$ condition is not reached inside BSs. We did not find the effect of scaling on the Love equations and on the moment of inertia in the literature. Hence, we show it here in Appendix. A and B.
The non-radial fundamental quasi-normal mode ($f$-mode) serves as a primary source of GW emission in compact stars. The relativistic Cowling approximation is often used to find mode frequency by neglecting metric perturbation; however, it neglects metric perturbation and overestimates the frequency as compared to that obtained within the framework of a general relativistic treatment.
In this study, the mode parameters are obtained by solving perturbations within the framework of linearized general relativistic treatment employing a direct integration method, as outlined in previous works [Detweiler85, Sotani2001, Pradhan2022], to solve the $f$-mode frequency of BSs. Numerical methods developed in a previous work [Pradhan2022] are employed for extracting the mode characteristics.
For the $f$-mode frequencies and damping times, we solve the $f$-mode eigenfrequencies using full-GR formalism. The details of the calculation of BS $f$-modes have been outlined in our other work [shirke_BS_fmodes_short]. There it was shown that if the $f$-mode frequency $f$ and damping time $\tau$ are scaled as:
$f$ $=f^{\prime}/(xM_{\rm Pl})$ (2.6)
$\tau$ $=\tau^{\prime}(xM_{\rm Pl})~,$ (2.7)
Then the $f$-mode equations, too, like the TOV equations, exhibit self-similarity, i.e., become independent of model parameters. The formalism and derivation of scaling have been provided in detail in the Appendix of [shirke_BS_fmodes_short].
@@ 3 Parameter Space
Fig. 1 shows the available parameter space in the $\lambda-m$ space for scalar DM. For mass, we show the range $10^{-25}\text{eV}<m<100\text{GeV}$ covering the mass of fuzzy DM or ultralight ALPs to roughly the mass of WIMPs. We consider $\lambda$ upto $\lambda=15$. This encompasses the upper limit of $\sim 4\pi$ as imposed by [Eby2016]. All the constraints are discussed in detail in the next section. The values $\lambda\lesssim 10^{-100}$ are irrelevant as lower values consistent with observations do not fall in the strong-interaction limit considered here. The yellow lines denote contours of fixed $\Lambda$, and this analysis is only applicable to $\Lambda>1000$, i.e., the strong-interaction regime marked by the yellow arrow. The region where $\Lambda\ll 1$ is the Kaup limit, and interactions can be ignored. This is shown just for depiction purposes. The results are not applicable in general below $\Lambda=1000$. The gray lines show contours where $\sigma/m$ (interaction cross-section per unit mass; see Eq. 3.1) is constant, and the bisque region between them is the region permitted by other astrophysical observations. We note that this is a very narrow region in the parameter space; however, the lower bound (dashed gray line) is not strict. The red line marks the constraint from CMB and LSS data [Cembranos2018]. The region above this line is excluded. The blue region is excluded by the recently set lower limit in scalar DM mass [Zimmerman2024] using dwarf galaxies.
@@
The maximum mass of the Bosonic dark star is shown in colour. This can be calculated for a given set of parameters $(\lambda,m)$ is given by [Colpi1986]: $M_{\rm max}=0.22\sqrt{\Lambda}\frac{M_{\rm Pl}^{2}}{m}=0.06\frac{\sqrt{\lambda}}{m^{2}}M_{\rm Pl}^{3}=0.06xM_{\rm Pl}^{3}$. The colour scale saturates into yellow at $M_{\rm max}=20^{22}M_{\odot}$, which is the rough estimate for the observable universe’s mass. We plot the contours of some fixed mass values. We find that the $M_{\rm max}$ ranges a larger region in mass spanning mountain mass to galactic mass (see caption of Fig. 1). The arrows indicate the regions allowed by various constraints. The only region bounded by the gray ($\sigma/m$), red (CMB, LSS), and blue (dwarf galaxies) lines is the allowed region for scalar DM and, hence, BSs. Further, this work is only applicable in the region bounded by the yellow $\Lambda=1000$ (strong-interaction regime) line.
@@
Flores et al. [Flores2019] restricted their study to stellar mass BSs and used selected values of $m\in\{1,1.25,1.5,1.75,2\}m_{n}$ and $a\in\{5,10,15,20\}$ fm to study the $f$-modes. The scattering length parameter $a=\lambda/(8\pi m)$ was used to fix the interaction strength instead of $\lambda$ in [Flores2019]. These are marked by a purple patch in Fig. 1. Maselli et al. [Maselli2017] studied the II-Love-$Q$ relations using a select set of model values. The parameters chosen were $m\in\{300,400\}$ MeV and $\lambda\in\{0,0.5,1.5\}\pi$. Celato et al. [Celato2025] explored the $f$-mode universal relations for the same parameters.  These are marked in orange in Fig. 1. It is clear that both these patches lie near the $M_{\rm max}\sim 1M_{\odot}$ contour and focus on the configurations in the NS mass range.
@@
Figure 1: Figure shows the available parameter space in the $\lambda-m$ space. The color indicates the logarithm of the maximum mass of a stable BS for a given pair of ($\lambda$,m). We show select contour lines (black) to get an idea of masses: $\log_{10}M=$ -18 (mountain mass; this is also the lower bound on primordial BHs set by Hawking radiation as lighter PBHs would evaporate within Hubble time); -9 (asteroid belt mass); -6 (Earth-like planet mass); -3 (giant planet mass); 0 (stellar mass); 2 (giant massive star mass); 6 and 9 (SMBH mass); 12 (galaxy mass). The yellow contours represent values of $\Lambda$. We restrict to the strong-interaction limit $\Lambda>1000$ in this work. All the results in this work, including the colorbar in this figure, are applicable only to this region. The orange and purple patch denotes the parameters used in earlier works on massive BSs in the strong-interaction regime [Maselli2017, Flores2019, Celato2025]. See text for details on various constraints.